The cell in an onion root tip cell takes 16 hours , calculate the length of time stage 2 lasts in a typical cell , give your answer to 2 significant figures
The table.
Stages in cell cycle
Non dividing cells s1 s2 s3 s4 total
Number of cells. 20. 9. 4. 2. 1. 36


S1 = stage one and so on
1
Expert's answer
2020-11-16T09:31:13-0500

Time for onion takes 16 hour

1 Hour = 60 minutes

Therefore 16 hours "\times" 60 minutes =960 minutes

Time for each stage=( Number of cells at each stage"\div" Total number of cells)"\times" 960


Time for stage two: Number of cells 9 cells

Total number of cell 72 cells


(9"\div"72 )"\times" 960 = 120 minutes

Stage 2 lasts for 120 minutes

Ans= 120 minutes